# Balanced Fibonacci word rectangles, and beyond

### v1

@@ Abstract
Following a recent paper of Anselmo et al., we consider $m\times n$ rectangular matrices formed from the Fibonacci word, and we show that their balance properties can be solved with a finite automaton. We also generalize the result to every Sturmian characteristic word corresponding to a quadratic irrational.
@@ 1 Introduction
Let $(a_{i})_{i\geq 0}$ be an infinite sequence over $\{0,1\}$. We can define an infinite matrix ${\cal A}=(a_{k,\ell})_{k,\ell\geq 0}$ by $a_{k,\ell}=a_{k+\ell}$. Then we can view the $m\times n$ submatrices of $\cal A$ as matrices where the successive rows form shifted length-$n$ “windows” of $(a_{i})_{i\geq 0}$, and the entry in the upper left corner has index sum $i$:
$A(i,m,n):=\begin{pmatrix}a_{i}&a_{i+1}&\dots&a_{i+n-1}\\ a_{i+1}&a_{i+2}&\dots&a_{i+n}\\ \vdots&\vdots&&\vdots\\ a_{i+m-1}&a_{i+m}&\dots&a_{i+m+n-2}\\ \end{pmatrix}$.
@@
It is clear that $A(i,m,n)=A(i,n,m)$ for all $i,m,n$.
The balance problem for $(a_{i})_{i\geq 0}$ is the following: for which pairs of natural numbers $(m,n)$ are there at most two distinct values among all the $(T(i,m,n))_{i\geq 0}$? If this is the case, we say that the $(m,n)$-submatrices of $A$ are balanced with respect to each other or just that $(m,n)$ is balanced. This is a generalization of a more familiar notion of balance for binary words, one that has been well-studied [15].
@@
for $n\geq 1$, where $\gamma=(3-\sqrt{5})/2$.
Recently Anselmo et al. [2] showed the following (their Theorem 18): for $m,n\geq 2$, if $\max(m,n)$ is a Fibonacci number, then the $m\times n$ blocks of ${\cal A}_{\gamma}$ are balanced. Also see [6]. However, there are still other $(m,n)$ pairs with the balance property, such as $(4,3)$. This raises the natural question, which $(m,n)$ pairs are balanced for the Fibonacci word?
@@ 2 Sturmian words
for $i\geq 1$. The infinite word $(a_{i})_{i\geq 1}$ is sometimes called a Sturmian characteristic word [7, 5]. Although by convention Sturmian words are indexed starting at $1$, we can define $a_{0}=0$ without creating any new blocks that do not already appear in $\bf f$, by a well-known result [5, 2.1.18].
We now define the infinite matrix ${\cal A}_{\alpha}$ from the terms of $(a_{i})_{i\geq 1}$; we call $A_{\alpha}$ a Sturmian matrix. By telescoping cancellation we see that
@@
The only thing left to show is that $(\Delta(i,m,n))_{i\geq 0}$ has a block of the form $1,0,0,\ldots,0,1$ or $1,0,0,\ldots,0,-1$ if and only if $(m,n)$ is not balanced.
@@ 3 Finite automata
Instead we can use the balance criterion stated in Lemma 1. This criterion is clearly expressible in first-order logic, since $x=\lfloor n\alpha\rfloor$ is expressible. We can then translate this first-order statement into an appropriate automaton that takes $m$ and $n$ as input (in the Ostrowski $\alpha$-numeration system) and accepts if and only if the $m\times n$ blocks of the matrix $A$ are balanced. We therefore get our main result.
@@
Let us now carry out this construction in Walnut for the Fibonacci word, where $\alpha=(3-\sqrt{5})/2$. We use the Zeckendorf numeration system [9, 16], which represents integers as a sum of distinct Fibonacci numbers $F_{j}$. A representation of $n$ is written compactly as a binary word $e_{i}\cdots e_{2}$, where $n=\sum_{2\leq j\leq i}e_{j}F_{j}$.
@@
The automaton in Figure 1 solves the balanced property for Fibonacci word rectangles.
@@
eval claima "?msd_fib Au,v,x,i,ip,j (\$adjfib(u,v) & x=u+v & i>v & ip>v
@@
eval claimb "?msd_fib Au,v,x,i,j (\$adjfib(u,v) & x=u+v & i>=1 & i<u & j>=1) =>
   (\$bal(v+i,j) <=> \$bal(x-i,j))":
@@ 4 Another version of the criterion for balance
We can obtain a more “human readable” version of the criterion for the $m\times n$ submatrices of $A_{\gamma}$ to be balanced, as follows:
@@ 5 Diverse rectangles
Consider the word rectangles formed from the Fibonacci word $\bf f$ with the indexing $f_{0}f_{1}f_{2}\cdots=0100$. The number of distinct values of $T(i,m,n)$, taken over all $i$, for $m=n=F_{3k}/2$, is at least $k+1$.

### v4

@@ Abstract
Following a recent paper of Anselmo et al., we consider $m\times n$ rectangular matrices formed from the Fibonacci word, and we show that their balance properties can be solved with a finite automaton. We also generalize the result to every Sturmian characteristic word corresponding to a quadratic irrational. Finally, we also examine the analogous question for the Tribonacci word and the Thue–Morse word.
@@ 1 Introduction
Let $(a_{i})_{i\geq 0}$ be an infinite sequence over $\mathbb{N}$. We can define an infinite matrix ${\cal A}=(a_{k,\ell})_{k,\ell\geq 0}$ by $a_{k,\ell}=a_{k+\ell}$. Then we can view the $m\times n$ submatrices of $\cal A$ as matrices where the successive rows form shifted length-$n$ “windows” of $(a_{i})_{i\geq 0}$, and the entry in the upper left corner has index sum $i$:
$A(i,m,n):=\begin{pmatrix}a_{i}&a_{i+1}&\dots&a_{i+n-1}\\ a_{i+1}&a_{i+2}&\dots&a_{i+n}\\ \vdots&\vdots&&\vdots\\ a_{i+m-1}&a_{i+m}&\dots&a_{i+m+n-2}\\ \end{pmatrix}$. (1)
@@
It is clear that $T(i,m,n)$ and $T(i,n,m)$ contain the same number of occurrences of each element of $(a_{i})_{i\geq 0}$, and so in particular $T(i,m,n)=T(i,n,m)$ for all $i,m,n$.
Now suppose that $(a_{i})_{i\geq 0}$ is an infinite sequence (or word) over $\{0,1\}$. In this case we call $A(i,m,n)$ a word rectangle. We say the $m\times n$ rectangles are balanced with respect to each other (or just that the pair $(m,n)$ is balanced ) if there are at most two distinct values among all the $(T(i,m,n))_{i\geq 0}$. This is a refinement of a more familiar notion of balance for binary words, one that has been well-studied [19].
From results of Berthé and Tijdeman [7] we know that $(m,n)$ is balanced for all $m,n$ only in trivial cases. Hence, the balance problem for $(a_{i})_{i\geq 0}$ is the following: determine which pairs $(m,n)$ of natural numbers are balanced.
@@
for $n\geq 1$, where $\gamma=(3-\sqrt{5})/2$. Recently Anselmo et al. [2] showed the following (their Theorem 18): for $m,n\geq 2$, if $\max(m,n)$ is a Fibonacci number, then the $m\times n$ word rectangles of the infinite Fibonacci matrix ${\cal A}_{\gamma}$ are balanced. Also see [7]. However, there are still other $(m,n)$ pairs with the balance property, such as $(4,3)$. This raises the natural question, which $(m,n)$ pairs are balanced for the Fibonacci word?
@@
We also examine the analogous question for the Tribonacci word and the Thue–Morse word.
@@ 2 Sturmian words
for $i\geq 1$. The infinite word $(a_{i})_{i\geq 1}$ is sometimes called a Sturmian characteristic word [8, 6]. Although by convention Sturmian words are indexed starting at $1$, we can define $a_{0}=0$ without creating any new blocks that do not already appear in $\bf f$, by a well-known result [6, 2.1.18].
We now define the infinite matrix ${\cal A}_{\alpha}$ from the terms of $(a_{i})_{i\geq 1}$; we call ${\cal A}_{\alpha}$ a Sturmian matrix. By telescoping cancellation we see that
@@
This fact can also be seen in another, perhaps simpler, way: since by Eq. (3) we know T(i+1,m,n)T(i,m,n)T(i+1,m,n)-T(i,m,n) is the difference in the number of $1$’s in the two length-$m$ words $a_{i+n}\cdots a_{i+m+n-1}$ and $a_{i}\cdots a_{i+m-1}$, Eq. (4) follows immediately from the fact that all Sturmian words are balanced.
@@
We claim that $(\Delta(i,m,n))_{i\geq 0}$ has a block of the form $1,0,0,\ldots,0,1$ or $1,0,0,\ldots,0,-1$ if and only if $(m,n)$ is not balanced.
@@ 3 Finite automata
Instead, we can use the balance criterion stated in Lemma 1. This criterion is clearly expressible in first-order logic, since $x=\lfloor n\alpha\rfloor$ is expressible. We can then translate this first-order statement into an appropriate automaton that takes $m$ and $n$ as input (in the Ostrowski $\alpha$-numeration system) and accepts if and only if the $m\times n$ blocks of the matrix $A$ are balanced. We therefore get our main result.
@@
Let us now carry out this construction in Walnut for the Fibonacci word, where $\alpha=\gamma=(3-\sqrt{5})/2$. We use the Zeckendorf numeration system [11, 20], which represents integers as a sum of distinct Fibonacci numbers $F_{j}$. A representation of $n$ is written compactly as a binary word $e_{i}\cdots e_{2}$, where $n=\sum_{2\leq j\leq i}e_{j}F_{j}$.
@@
The automaton in Figure 1 solves the balance property for the Fibonacci word rectangles of dimension $m\times n$.
@@
eval claim_a "?msd_fib Au,v,x,i,ip,j (\$adjfib(u,v) & x=u+v & i>v & ip>v
@@
eval claim_b "?msd_fib Au,v,x,i,j (\$adjfib(u,v) & x=u+v & i>=1 & i<u & j>=1)
   => (\$bal(v+i,j) <=> \$bal(x-i,j))":
@@ 4 Another version of the criterion for balance
We can obtain a more “human readable” version of the criterion for the $m\times n$ submatrices of ${\cal A}_{\gamma}$ to be balanced, as follows:
@@ 5 Diverse rectangles
Consider the word rectangles formed from the Fibonacci word $\bf f$ with the indexing $f_{0}f_{1}f_{2}\cdots=0100\cdots$. The number of distinct values of $T(i,m,n)$, taken over all $i$, for $m=n=F_{3k}/2$, is at least $k+1$.
@@
6 Balanced rectangles more generally
Let us look at the notion of balanced rectangles somewhat more generally. Given a word $w$ and a symbol $c$, we let $|w|_{c}$ denote the number of occurrences of $c$ in the word $w$. We say that an infinite sequence $\bf a$ is $r$-balanced if, for all $c$ and all pairs of equal-length blocks $x$ and $y$ occurring in $\bf a$, we have $\left||x|_{c}-|y|_{c}\right|\leq r$.
We can generalize this to word rectangles in the obvious way: if $A(i,m,n)$ is defined as in Eq. (1), then $|A(i,m,n)|_{c}$ is the total number of occurrences of the letter $c$ in the $m\times n$ rectangle $A(i,m,n)$. Then the natural question is, for which pairs $(m,n)$ are all the rectangles $A(i,m,n)$ $r$-balanced with respect to each other? That is, for which $m,n$ do we have $\left||A(i,m,n)|_{c}-|A(j,m,n)|_{c}\right|\leq r$ for all $i,j$ and all $c$?
Consider the infinite Tribonacci word ${\bf TR}=0102010\cdots=t_{0}t_{1}t_{2}\cdots$, the fixed point of the morphism $0\rightarrow 01$, $1\rightarrow 02$, and $2\rightarrow 0$; see, for example, [4]. It seems natural to examine the analogous question of balance for the word rectangles formed from this word. It is known (see [14]) that the Tribonacci word is $2$-balanced; that is, for all $c\in\{0,1,2\}$ and $n\geq 0$, the number of occurrences of the letter $c$ in two words of length $n$ differs by at most $2$. Here is the problem we solve in this section: characterize $m,n$ such that all the $(m,n)$-rectangles built from the Tribonacci word are $2$-balanced with respect to each other. By symmetry, it suffices to answer this question for $m\leq n$.
Our proof technique is, once again, based on automata. Here we need automata that take the Tribonacci representation [10] of an integer as input. However, the technique is somewhat different than the one we used for Fibonacci word rectangles.
Theorem 10.
Assume $m\leq n$. Then
  • (a)
    If $m=1$, then all the $m\times n$ rectangles are $2$-balanced;
  • (b)
    If $m=2$, there is a $77$-state Tribonacci automaton that accepts exactly those $n$ for which all the $m\times n$ rectangles are $2$-balanced;
  • (c)
    If $m\geq 3$, then there are no $n$ for which the all the $m\times n$ rectangles are $2$-balanced.
Proof.
  • (a)
    Trivially follows from the $2$-balance of the Tribonacci word.
  • (b)
    We use Walnut to create the automaton. The code for triba, tribb, etc. is taken from [17, §10.12].
    reg shift {0,1} {0,1} "([0,0]|[0,1][1,1]*[1,0])*":
    def triba "?msd_trib (s=0&n=0) | Ex \$shift(n-1,x) & s=x+1":
    # position of the n’th 0 in TR, indexing starting with 1
    def tribb "?msd_trib (s=0&n=0) | Ex,y \$shift(n-1,x) &
       \$shift(x,y) & s=y+2":
    # position of the n’th 1 in TR, indexing
    starting with 1
    def tribc "?msd_trib (s=0&n=0) | Ex,y,z \$shift(n-1,x) &
       \$shift(x,y) & \$shift(y,z) & s=z+4":
    # position of the n’th 2 in TR, indexing
    starting with 1
    def tribd "?msd_trib Et,u \$triba(s,t) & \$triba(s+1,u) & t<=n & n<u":
    # s = number of 0’s in TR[0..n-1]
    def tribe "?msd_trib Et,u \$tribb(s,t) & \$tribb(s+1,u) & t<=n & n<u":
    # s = number of 1’s in TR[0..n-1]
    def tribf "?msd_trib Et,u \$tribc(s,t) & \$tribc(s+1,u) & t<=n & n<u":
    # s = number of 2’s in TR[0..n-1]
    def num0 "?msd_trib Ex,y \$tribd(i,x) & \$tribd(i+n,y) & z+x=y":
    # z is the number of 0’s in TR[i..i+n-1]
    def num1 "?msd_trib Ex,y \$tribe(i,x) & \$tribe(i+n,y) & z+x=y":
    # z is the number of 1’s in TR[i..i+n-1]
    def num2 "?msd_trib Ex,y \$tribf(i,x) & \$tribf(i+n,y) & z+x=y":
    # z is the number of 2’s in TR[i..i+n-1]
    # what follows is the code that generates the automaton accepting the n
    # such that the 2xn rectangles are 2-balanced
    def row2_0 "?msd_trib Ex,y \$num0(i,n,x) & \$num0(i+1,n,y) & z=x+y":
    # z is number of 0’s in 2xn rectangle cornered at i
    def row2_1 "?msd_trib Ex,y \$num1(i,n,x) & \$num1(i+1,n,y) & z=x+y":
    # z is number of 1’s in 2xn rectangle cornered at i
    def row2_2 "?msd_trib Ex,y \$num2(i,n,x) & \$num2(i+1,n,y) & z=x+y":
    # z is number of 2’s in 2xn rectangle cornered at i
    def chk2_0 "?msd_trib Ai,j,y,z (\$row2_0(i,n,y) & \$row2_0(j,n,z) & y<=z)
       => (z<=y+2)":
    # check 2-balance of 0’s in 2xn matrices
    def chk2_1 "?msd_trib Ai,j,y,z (\$row2_1(i,n,y) & \$row2_1(j,n,z) & y<=z)
       => (z<=y+2)":
    # check 2-balance of 1’s in 2xn matrices
    def chk2_2 "?msd_trib Ai,j,y,z (\$row2_2(i,n,y) & \$row2_2(j,n,z) & y<=z)
       => (z<=y+2)":
    # check 2-balance of 2’s in 2xn matrices
    def bal2 "?msd_trib \$chk2_0(n) & \$chk2_1(n) & \$chk2_2(n)":
    # 77 states
    The first few such $n$ are
    $1,2,3,4,7,8,9,10,11,14,15,22,23,24,27,28,33,34,35,46,47,48,\ldots.$
    The fact that we need 77 states suggests there is unlikely to be a simple English description of these $n$.
  • (c)
    Once again we use Walnut, but with a different idea. Define the matrix $B(i,m,n)$ to be $A(i,m,n)$ with the letters recoded as follows: $0,1\rightarrow 0$ and $2\rightarrow 2$. The idea is to show that provided $m,n\geq 3$ there exist $i,j$ such that $B(i,m,n)$ and $B(j,m,n)$ agree, except on the $3\times 3$ submatrix lying in the lower right corrner: $B(i,m,n)$ has
    $\left[\begin{matrix}0&0&2\\ 0&2&0\\ 2&0&0\end{matrix}\right]$
    in the lower right corner, while $B(j,m,n)$ has
    $\left[\begin{matrix}0&0&0\\ 0&0&0\\ 0&0&0\end{matrix}\right].$
    Thus $|A(i,m,n)|_{2}\geq|A(j,m,n)|_{2}+3$, showing that the $m\times n$ rectangles cannot be $2$-balanced with respect to each other.
    We use the following Walnut code:
    morphism map2 "0->0 1->0 2->2":
    image TR2 map2 TR:
    def trib2eqfac "?msd_trib Au,v (u>=i & u<i+n & u+j=v+i) =>
       TR2[u]=TR2[v]":
    eval cb2 "?msd_trib Ap Ei,j \$trib2eqfac(i,j,p) & TR2[i+p]=@0 &
       TR2[i+p+1]=@0 & TR2[i+p+2]=@2 & TR2[i+p+3]=@0 & TR2[i+p+4]=@0 &
       TR2[j+p]=@0 & TR2[j+p+1]=@0 & TR2[j+p+2]=@0 & TR2[j+p+3]=@0 &
       TR2[j+p+4]=@0":
    And Walnut returns TRUE.
    Here TR2 is an automaton for the Tribonacci word ${\bf TR}_{2}$ that is 0 at positions where the Tribonacci word has values 0 or 1, and 2 at positions with value 2. The automaton trib2eqfac accepts $(i,j,n)$ if and only if ${\bf TR}_{2}[i..i+n-1]={\bf TR}_{2}[j..j+n-1]$.
    Finally, cb2 says that for all $p$ there exist $i,j$ such that the block of length $p+5$ starting at $i$ looks like $w00000$, while the block of length $p+5$ starting at $j$ looks like $w00200$. Thus, if $m,n\geq 3$, we can set p=m+n6p=m+n-6 to find two $m\times n$ rectangles that are unbalanced with respect to each other for the symbol $2$.
7 The Thue–Morse word
We can also form rectangles out of the Thue–Morse word ${\bf t}=01101001\cdots$, fixed point of the morphism $0\rightarrow 01$, $1\rightarrow 10$. Here we can get a finite automaton for a complete description of the balance. It is easier in this case because $t_{2n+1}=1-t_{2n}$, so one can “pair up” entries except at the edges of the rectangle.
Theorem 11.
  • (a)
    Every $m\times n$ word rectangle $A(i,m,n)$ built from $\bf t$ satisfies $|2\cdot|A(i,m,n)|_{1}-mn|\leq 4$.
  • (b)
    If there is an $m\times n$ word rectangle with $2|A(i,m,n)|_{1}-mn=c$, then there is one with $2|A(j,m,n)|_{1}-mn=-c$.
  • (c)
    The $m\times n$ word rectangles, for $m,n\geq 1$ have balance either $1,2,3$ or $4$.
Proof.
Let us write ${\bf t}=01101001\cdots=t_{0}t_{1}t_{2}\cdots$. (This is the same notation we used for Tribonacci but hopefully the reader will not be confused.) It follows from an easy induction that $t_{2k}+t_{2k+1}=1$ for all $k\geq 0$. Similarly, $t_{4k}+t_{4k+2}=1$ and $t_{4k+1}+t_{4k+3}=1$ for all $k\geq 0$. Recall the definition (1) of $A(i,m,n)$. Since each row contains a block of $\bf t$}, we can use the fact $t_{2k}+t_{2k+1}=1$ to pair up adjacent bits with sum $1$ in each row of $A(i,m,n)$. Then exactly the following bits will be left unmatched: all entries in the leftmost column with odd indices and all entries in the rightmost column with even indices. In the second step we can use $t_{4k+1}+t_{4k+3}=1$ to pair up bits in the leftmost column, and use $t_{4k}+t_{4k+2}=1$ to pair up bits in the rightmost column. At the end, at most four bits might be left unmatched, namely
  • either $a_{i}$ or $a_{i+1}$, in case one of them is congruent to $3$ modulo $4$;
  • either $a_{i+m-2}$ or $a_{i+m-1}$, in case one of them is congruent to $1$ modulo $4$;
  • either $a_{i+n-1}$ or $a_{i+n}$, in case one of them is congruent to $2$ modulo $4$;
  • either $a_{i+n+m-3}$ or $a_{i+n+m-2}$, in case one of them is congruent to 0 modulo $4$.
Now note that we have matched at least mn4mn-4 bits, whose sum is exactly (mn4)/2=mn/22(mn-4)/2=mn/2-2. This immediately implies $mn/2-2\leq T(i,m,n)\leq mn/2-2+4=mn/2+2$. This gives the bound in (a).
(b) follows because if there is a word rectangle with $2|A(i,m,n)|_{1}-mn=c$, then there is a word rectangle with every 0 flipped to 1 and every 1 flipped to 0, giving a difference of c-c.
(c) follows from (a) and the fact that the Thue–Morse sequence is not eventually periodic, so we cannot have balance 0. ∎
Now let us create a Walnut automaton to compute the balance. This introduces some complications. Possibly the conceptually simplest way involves computation with negative numbers, and the simplest way to do that in Walnut is to represent integers in base-(2)(-2). Then, at the end, when getting the automaton for balance, we can convert it to a base-$2$ automaton using the techniques of [18].
Theorem 12.
There is an automaton of $92$ states that on input $m$ and $n$ in base $2$ determines the balance of the $m\times n$ rectangles built from $\bf t$.
Proof.
We use the following Walnut code. The idea is illustrated for the case $i,m,n$ all even. We consider pairing all the entries of the matrix with an even index with the odd index that follows it. This leaves unpaired entries in the leftmost column and the rightmost column.
In the leftmost column these entries are $t_{i+1}+t_{i+3}+\cdots+t_{i+m-1}$. Since all the indices are odd, this is the same as $(1-t_{i/2})+(1-t_{(i+2)/2})+\cdots+(1-t_{(i+m-2)/2})=m/2-(t_{i/2}+t_{i/2+1}+\cdots+t_{(i+m)/2-1})$. In the rightmost column these entries are $t_{i+n}+t_{i+n+2}+\cdots+t_{i+m+n-2}=t_{(i+n)/2}+t_{(i+n)/2+1}+\cdots+t_{(i+m+n)/2-1}$. The case of $i$ odd is similar, and we reduce the other cases to this one by removing a row or column, or both.
Here the automaton TM2 computes the Thue-Morse sequence, but with input given in base-(2)(-2). The automaton twonegtwo takes two inputs, $w$ and $x$ in parallel, and accepts if $w$ considered as a base-(2)(-2) expansion equals $x$ considered as a base-$2$ expansion. Both can be downloaded from the website of the first author.
def n2even "?msd_neg_2 Ek n=2*k":
def n2odd "?msd_neg_2 Ek n=2*k+1":
def prefsum "?msd_neg_2 (s=n/2 & \$n2even(n)) | (s=n/2+TM2[n-1] & \$n2odd(n))":
# sum of t[0..n-1]
# 17 states
def factorsum "?msd_neg_2 Eu,v \$prefsum(i,u) & \$prefsum(i+n,v) & s+u=v":
# sum of t[i..i+n-1]
# 317 states
def m_even_n_even "?msd_neg_2 Ea,b \$factorsum(i/2,m/2,a) &
   \$factorsum((i+n)/2,m/2,b) & s=2*(b-a)":
# s = 2T(i,m,n) - mn  for m even, n even
def m_even_n_odd "?msd_neg_2 Ea,b \$m_even_n_even(i,m,n-1,a) &
   \$factorsum(i+n-1,m,b) & s=a+2*b-m":
def m_odd_n_even "?msd_neg_2 Ea,b \$m_even_n_even(i,m-1,n,a) &
   \$factorsum(i+m-1,n,b) & s=a+2*b-n":
def m_odd_n_odd "?msd_neg_2 Ea,b \$m_odd_n_even(i,m,n-1,a) &
   \$factorsum(i+n-1,m,b) & s=a+2*b-m":
def tmt "?msd_neg_2  i>=0 & m>=0 & n>=0 &
((\$n2even(m) & \$n2even(n) & \$m_even_n_even(i,m,n,s)) |
(\$n2even(m) & \$n2odd(n) & \$m_even_n_odd(i,m,n,s)) |
(\$n2odd(m) & \$n2even(n) & \$m_odd_n_even(i,m,n,s)) |
(\$n2odd(m) & \$n2odd(n) & \$m_odd_n_odd(i,m,n,s)))"::
# computes 2T(i,m,n) - mn for all i, m, n >= 0
# 3678 states
def maxval "?msd_neg_2 (Ei \$tmt(i,m,n,x)) & Ai,s \$tmt(i,m,n,s) => s<=x":
def tmbal "Ex,y,b \$twonegtwo((?msd_neg_2 x),m) & \$twonegtwo((?msd_neg_2 y),n) &
   \$twonegtwo((?msd_neg_2 b),z) & \$maxval(?msd_neg_2 x,y,b)"::
def aut0 "\$tmbal(m,n,0)":
def aut1 "\$tmbal(m,n,1)":
def aut2 "\$tmbal(m,n,2)":
def aut3 "\$tmbal(m,n,3)":
def aut4 "\$tmbal(m,n,4)":
combine TMB aut0=0 aut1=1 aut2=2 aut3=3 aut4=4:
# result has 92 states
Theorem 13.
Suppose $m,n\geq 3$. The set of $m\times n$ word rectangles for $\bf t$ has balance exactly $3$ if and only if $m$ and $n$ are both odd. Here by “balance exactly 33” we mean the balance is bounded above by $3$, and there is also a pair of rectangles that achieves this bound.
Proof.
We use the following Walnut code:
def even "Ek n=2*k":
def odd "Ek n=2*k+1":
eval bal3 "Am,n (m>=3 & n>=3) => ((\$odd(m) & \$odd(n)) <=> TMB[m][n]=@3)":
And Walnut returns TRUE. ∎
The $m\times n$ rectangles having balance exactly $2$ or $4$ are more difficult to understand.